Fermi National Laboratory


Accelerator Update for May 16, 2005 - May 18, 2005

Monday May 16
The day shift began with Operations monitoring store 4147, the Recycler stash, and the Antiproton Source (Pbar) stack, and with all beam beyond Booster off due to problems with the Booster chopper.

At 8:55 AM, a Booster RF station (BRF16) tripped off. It reset okay.

Booster experts reported at 9:19 AM that they had repaired the chopper and that they were tuning Booster.

Operations established beam to Pbar, and to the MiniBooNE, NuMI, and the Meson MCenter experiments at 9:20 AM.

At 10:34 AM, BRF16 tripped off. It reset okay.

At 11:46 AM, BRF12 tripped off many times. At 12:03 PM, it tripped off again. It tripped on a ground fault indication. Experts investigated.

At 12:05 PM, the Booster East Anode Power Supply (EAPS) tripped off. Eventually, it reset.

At 12:40 PM, EAPS tripped off twice more. Experts were there and reset the supply. It tripped off again four minutes later.

At 1:22 PM, Operations reported that the beam permit for NuMI dropped out due to the voltage on a power supply (HV101) dipping low, but a plot showed that the power supply was at the proper level. Operators contacted an expert.

At 2:44 PM, BR12 tripped off. Operators investigated. They reported that it reset okay, but wouldn't stay on for more that a couple of minutes. Experts reported that the problem was in the tunnel. Operators left the station off.

At 6:16 PM, Operations reported that the EAPS tripped off for the first of two times in fifteen minutes. It reset okay.

Operations terminated store 4147 at 9:02 PM.

Operations put the Tevatron (TeV) into shot setup at 9:23 PM.

At 9:46 PM, the Booster chopper tripped off.

Operations established combination store 4148 at 11:38 PM with an initial luminosity of 111.45E30. Pbar resumed stacking and NuMI resumed taking beam.

MCenter resumed taking beam at 11:49 PM.

Tuesday May 17
The midnight shift began with Operations monitoring store 4148, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations reported at 12:05 AM that the I- Source had sparked 150 times during the last eight hours. Much later, Preaccelerator experts reported that the sparking was due to a problem with the water resistor.

At 3:33 AM, BRF16 tripped off. It reset okay.

At 4:33 AM, BRF17 tripped off for the first of two times in four minutes. It reset okay both times.

Operators turned off beam in the Main Injector at 6 AM in preparation for an access later this morning.

At 8:59 AM, Pbar emittances blew up. Operators worked at fixing the problem.

MiniBooNE experts reported at 10:52 AM that they were trying to start their foil studies, but beam problems were delaying them. They finally finished the study at 11:11 AM.

Operations shut off all beam at 12:01 PM for work in Linac, Booster, Pbar, and MI.

At 3:02 PM, technicians reported a LCW leak on the MI dump. Operators notified the water techs and the techs had the leak repaired within forty-five minutes.

Operations reported that both Linac and Booster were back on by 3:17 PM.

Operations reported at 4:08 PM that MI was ramping.

At 4:10 PM, Operations reported trouble with a Linac RF station (LRF1). An operator notified an expert.

Pbar resumed stacking at 4:21 PM.

At 4:32 PM, BRF12 tripped off for the first of several times over the next hour.

Operations established beam to MiniBooNE at 5:22 PM.

At 6:21 PM, the Booster chopper tripped off. It would not stay reset. An operator called in experts. Eventually, the experts found that a cooling fan had seized up, which caused a cable to overheat. While making repairs they noticed that another fan was failing.

The Booster experts reported at 9:51 PM that the chopper was up and running.

Pbar resumed stacking at 9:52 PM.

At 9:52 PM, BRF12 tripped off on a ground fault indication. It tripped off two more times in the next two minutes. Operations left the station off after the last trip.

Operations put the Recycler into shot setup to stash antiprotons at 10:31 PM. They finished stashing an hour later. NuMI took beam for the first time during a shot setup. This was a test.

Pbar resumed staking at 11:30 PM.

Wednesday May 18
The midnight shift began with Operations monitoring store 4148 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

Operations reestablished beam to MCenter at 12:09 AM.

At 12:29 AM, MIRF18 tripped off on a LCW flow fault. After two minutes, operators were able to reset the station.

At 2:39 AM, BRF16 tripped off. It reset okay.

Operations terminated store 4148 at 6:03 AM. Unfortunately, the TeV suffered a sector F4 quench right after the termination.

Operators made the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls ready for access at 6:33 AM.

At 7:50 AM, Pbar stopped stacking due to the failure of a controls card. An expert investigated and began repairs.

The Plan for Wednesday
The plans for today are to repair Pbar, allow CDF and Do to finish their access, allow Booster experts to do more chopper maintenance, and then establish a new store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/18/2005   email Fermilab